OpenStudy (anonymous):

10m^4/9a^2 times 3a^2b

OpenStudy (anonymous):

\[ \frac {10m^4}{9a^2} 3a^2 b \]

OpenStudy (anonymous):

\[ \frac {10m^4}{9a^2} 3a^2 b=\frac {10m^4}{3} b=\frac {10 b m^4}{3} \]